A vertical wire with a diameter of 3.60 mm and an unstretched length of 4.25 m extends by a distance of 9.19 mm when it is loaded with a 60N weight.

Calculate the strain in the wire


1
Expert's answer
2021-01-28T19:39:07-0500

By the definition of the strain, we get:


"\epsilon=\dfrac{\Delta L}{L_0}=\dfrac{L-L0}{L_0},""\epsilon=\dfrac{4.25919\ m-4.25\ m}{4.25\ m}=2.16\cdot10^{-3}."